OpenStudy (anonymous):

how to graph the equation y = 4(2)x

OpenStudy (anonymous):

its in y=mx +b form so what is the y intercept and the slope

OpenStudy (anonymous):

In the equation y = mx + b, m is the slope and b is the y-intercept, I think.

OpenStudy (anonymous):

yes so m = 4(2) = 8 and b = 0